What is the deduce of the following physical quantities:
(1) Frequency
(2) Kinetics

Grade:10

1 Answers

Harshit Singh
3 years ago
Dear Student

1. Frequency = 1/ time period = 1/ T = T-1
2.Kinetics: It’s not a quantity maybe you have done typing error.

Thanks